A CORRECTION TO “EXPANSIONS OF O-MINIMAL STRUCTURES BY FAST SEQUENCES”

CHRIS MILLER

Version: December 1, 2018

This is not a preprint; please do not refer to it as such.

It has been noticed (by A. Dolich) that the second sentence of [1, 2.1] is easily seen to be false; it should simply be deleted because it is never actually used. What is true: If A is homogeneous and f is indiscernible, then y ↦ f(x, y) is indiscernible; and if f is almost indiscernible, then y ↦ f(x, y) is almost indiscernible.

References

[1] Harvey Friedman and Chris Miller, Expansions of o-minimal structures by fast sequences, J. Symbolic Logic 70 (2005), no. 2, 410–418, DOI 10.2178/jsl/1120224720. MR 2140038

Department of Mathematics, The Ohio State University, 231 West 18th Avenue, Columbus, Ohio 43210, USA

  E-mail address: miller@math.osu.edu